Consider a consumer who is choosing how many of two goods to buy: Footballs and cricket balls. The consumer has an income of $20, and the cost of a football is $4 and a cricket ball is $2. (a) Write down the equation for the consumer’s budget constraint and graph it.(b) The government decides that football is evil and needs to be taxed. They introduce a 50% tax on each football sold. Rewrite and re-graph the budget constraint.(c) A new government is elected that hates all sports. They now tax both footballs and cricket balls at 50%. What does the budget constraint look like now?
1
Expert's answer
2021-08-06T08:12:09-0400

a.

Y=20

f= 4

c=2

Budget constraint is given by; "y= 4f + 2c"

Hence; "BC= 4F+2C = 20"











b.

New cost of football = 1.5 x 4 = 6

New budget constraint is


"6f + 2c = 20"











c.

The costs football and cricket will increase by 50%

New costs will be;


"football = 1.50 x 4 = \$6"


"cricket = 1.50 x 2 = \$ 3"

Therefore the new budget constraint will be


"6f + 3c = 20"

The new budget constraint will be more flatter, meaning more crickets will be bought.